OpenStudy (anonymous):

write the fraction as a mixed number in simplest form 42/4 and 31/12

jimthompson5910 (jim_thompson5910):

42/4 = (40+2)/4 = 40/4 + 2/4 = 10 + 1/2 = 10 1/2 So 42/4 = 10 1/2 I'll let you do the other one

jimthompson5910 (jim_thompson5910):

The key is to break up the numerator so that one of the pieces is a multiple of the denominator (as big as you can get)

OpenStudy (anonymous):

No! im going to be wronhe

OpenStudy (anonymous):

wrong*

OpenStudy (anonymous):

Hold on lemme try one more time :d

jimthompson5910 (jim_thompson5910):

just give it a shot and show me what you get, if it's wrong, then I'll help you out with the second

OpenStudy (anonymous):

1 3/4?

jimthompson5910 (jim_thompson5910):

how did you get that?

OpenStudy (anonymous):

how many times can 7 go into 8 1 time with 3/4 left over

jimthompson5910 (jim_thompson5910):

the problem is 31/12 right?

jimthompson5910 (jim_thompson5910):

If so, then 31/12 = (24 + 7)/12 = 24/12 + 7/12 = 2 + 7/12 = 2 7/12 This means that 31/12 = 2 7/12

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

oh.. :P sorry

jimthompson5910 (jim_thompson5910):

Alternative method: Using a calculator, 31/12 = 2.583333 So the whole part is 2 and the fractional part is 0.583333 Ignore the decimal fractional part Since the whole part is 2, we know that 12*2 = 24 is less than 31 Now subtract this from 31 to get 31 - 24 = 7 This is the numerator that will go over the 12 So 31/12 = 2 7/12